### v2

@@ 3.3 Summary Statistics
Table 1: Summary statistics of the mid-quote return, order flow imbalance, number of order book events, average size of order book events, average spread, and depth defined in (1). The statistics are computed from 34,512,298 samples of returns, flow, and the number and average size of events, and from 5,874,016 samples of the depth, computed for one-second intervals.
@@
in basis points, ∗∗ in the hundreds, ∗∗∗ in the thousands
@@ 3.4 Intraday Variations
Figure 1: Standard deviations of the mid-quote returns and order flow imbalances and means of the number of events, average size of events, average spread, and depth. The statistics are computed for each five-minute interval from 8:30 until 15:00. The dashed lines indicate the statistics calculated from all samples.
@@ 4.1 Simple Bivariate Model
This implies $y_{t}=B^{-1}\epsilon_{t}$, where $B^{-1}$ is non-diagonal, meaning $f_{t}$ and $\epsilon_{r,t}$ are correlated. As a result, OLS applied to (2) produces an inconsistent estimator of $b_{r}$ when $b_{f}\neq 0$. To obtain consistent estimates, we adopt the ITH method described below.
@@ 4.2 ITH Method
From (4), the relationship $B\Sigma B^{\prime}=\Omega$ holds, which yields three equations:
@@
The ITH method resolves this by exploiting heteroskedasticity. Suppose there are $S$ states such that the variance–covariance matrix $\Sigma$ changes across states, while $b_{r}$ and $b_{f}$ in $B$ remain constant. For each state $s=1,\dots,S$, equations (7)–(9) hold, with $\omega_{r,s}$ and $\omega_{f,s}$ varying by state.
@@
the parameters are identified. [^8] See Rigobon (2003) for details on estimation and identification in general settings. Under these conditions, the parameters can be estimated by the generalized method of moments (GMM) using the $3S$ moment conditions implied by (7)–(9).
@@ 4.3 SVAR Model
where $B$ and $\Omega$ are as in (5) and $\Phi_{j}$ are $2\times 2$ matrices capturing autocorrelation.
@@ 5.1 Macroeconomic News Effects
Figure 2: Average values of structural parameters ($b_{r}$, $b_{f}$, $\omega_{r}$, and $\omega_{f}$) and market activity variables ($D^{-1}$ and $SPR$) on days with macroeconomic announcements released at 9:00 Chicago Time (black circles, solid line) and on days without announcements (gray triangles, dashed line). Estimates of $b_{r}$ and $b_{f}$ are averaged over 15-minute intervals, while the other variables are averaged over five-minute intervals.
To formalize these patterns, we estimate regressions with two announcement dummies. The first, $ANN_{t}$, equals one if any of the eight announcements are released in interval $t$; the second, $ANN_{t}^{-}$, equals one if the announced value falls below the consensus forecast. Leads and lags of these dummies are included to capture pre- and post-announcement effects. Regressions are run separately for parameters computed over 15-minute intervals ($b_{r}$ and $b_{f}$) and over 5-minute intervals ($\omega_{r}$ and $\omega_{f}$), with controls for market activity variables—the reciprocal of depth $D^{-1}$ defined in (1), the number of order book events $NE$, the average size of events $ASE$, and the average spread SPRSPR—as well as time dummies.
@@
Table 2: Regression results with the announcement dummies. The dependent variables are the price impact ($b_{r}$), flow impact ($b_{f}$), return innovation volatility ($\omega_{r}$), and flow innovation volatility ($\omega_{f}$). The coefficients of the announcement dummy $ANN_{t}$, negative-announcement dummy $ANN_{t}^{-}$, and their leads and lags, if significant, are reported. The coefficients of the other independent variables are almost identical to those in Table 5 and are omitted for brevity. $\bar{R}^{2}$ indicates the adjusted R-squared value. Robust standard errors adjusted for 1,490 clusters in date in parentheses.
@@
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@ 5.2 Structural Parameters
The reduced-form VAR model (12) is estimated using returns $r_{t}$ and order flow imbalances $f_{t}$ over one-second intervals, separately for each 15-minute interval and each day. This results in 38,740 estimations. [^10] The number of lags $p$ chosen by the Akaike information criterion is positive in 38,725 intervals. On average, lagged returns explain 8.7% of the variation in returns and lagged flows explain 4.6% of the variation in flows, with $R^{2}$ values occasionally exceeding 20%. These findings confirm strong auto- and cross-correlation structures, motivating the SVAR framework. The variance–covariance matrices of the residuals are then used to check the rank condition (10), leaving 37,029 intervals suitable for ITH estimation.
For each valid interval, the ITH method is applied to estimate the structural parameters in (5). The price and flow impacts ($b_{r},b_{f}$) are assumed constant within a 15-minute window, while the standard deviations of return and flow innovations ($\omega_{r},\omega_{f}$) vary across three nested five-minute subintervals. This yields one set of $(b_{r},b_{f})$ and three sets of $(\omega_{r},\omega_{f})$ for each 15-minute block. Because three heteroskedastic states are used, the system is over-identified.
@@
Table 3: Estimation results of the structural parameters in (5). For each 15-minute interval satisfying the rank condition (10), the price and flow impacts $(b_{r},b_{f})$ and three sets of standard deviations $(\omega_{r,1},\omega_{r,2},\omega_{r,3},\omega_{f,1},\omega_{f,2},\omega_{f,3})$ are estimated by using the ITH method, resulting in 37,029 estimations. The last column * represents the proportion of intervals where the coefficient is statistically significant, that is, the $t$-value is greater than two in absolute value.
@@ 5.3 Impulse Responses
Return-to-Return Impact $IRF_{rr}$ Flow-to-Return Impact $IRF_{rf}$
@@
Return-to-Flow Impact $IRF_{fr}$ Flow-to-Flow Impact $IRF_{ff}$
@@
Figure 3: Impulse responses induced by a one-unit shock in return and flow innovations up to 10 lags. $IRF_{ij}$ represents the response of $i$ to the shock in $j$. The medians of the 37,029 estimates are presented. The shaded area indicates the 5th and 95th percentiles.
Table 4 reports summary statistics of the long-run impacts $I_{ij}(\infty)$ defined in (13). All are positive on average except for $I_{fr}(\infty)$ in some intervals. A one-unit flow shock, corresponding to an unexpected increase of about 1,000 contracts at the best bid or ask, raises prices by roughly 1.42 basis points in the long run. Using the mean flow volatility $\omega_{f}$ (0.34, about half the mean depth in Table 1), this translates into a price impact of 0.48 basis points—closely aligned with the 0.5–0.6 basis points reported by Hautsch and Huang (2012) for limit orders of half-depth.
@@
Table 4: Summary statistics of the 37,029 estimates of the long-run impacts $I_{ij}(\infty)$ defined in (13).
@@ 5.4 Intraday Variations
Figure 4: Intraday averages of estimated price and flow impacts ($b_{r},b_{f}$) and standard deviations of return and flow innovations ($\omega_{r},\omega_{f}$). The averages of $b_{r}$ and $b_{f}$ are computed for each 15-minute interval; those of $\omega_{r}$ and $\omega_{f}$ are computed for each five-minute interval. Dashed lines indicate full-sample averages.
@@
Table 5: Regression results for intraday variations. Dependent variables are structural parameters ($b_{r},b_{f}$) estimated over 15-minute intervals and innovation volatilities ($\omega_{r},\omega_{f}$) estimated over five-minute intervals. Explanatory variables are reciprocal depth $D^{-1}$ (per thousand), number of events NENE (in millions), average size of events ASEASE (in thousands), and average spread SPRSPR, plus time dummies. Robust standard errors are clustered by date.
@@
Robust standard errors adjusted for 1,490 date clusters.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$

### v4

@@ 3.3 Summary Statistics
Table 1: Summary statistics of mid-quote returns, OFI, and market activity variables
@@
  • Notes: Statistics are computed from 34,512,298 samples of returns, flow, and the number and average size of events, and from 5,874,016 samples of depth, all measured at one-second intervals. Depth is defined in Eq. (1). in basis points, ∗∗ in hundreds, and ∗∗∗ in thousands.
@@ 3.4 Intraday Variations
Figure 1: Standard deviations of mid-quote returns and order flow imbalances, and means of the number of events, average event size, average spread, and depth. Statistics are computed for each five-minute interval between 8:30 and 15:00. Dashed lines indicate the overall averages across all samples.
@@ 4.1 Simple Bivariate Model
This implies $y_{t}=B^{-1}\epsilon_{t}$, where $B^{-1}$ is non-diagonal, meaning $f_{t}$ and $\epsilon_{r,t}$ are correlated. As a result, OLS applied to Eq. (2) produces an inconsistent estimator of $b_{r}$ when $b_{f}\neq 0$. To obtain consistent estimates, we adopt the ITH method described below.
@@ 4.2 ITH Method
From Eq. (4), the relationship $B\Sigma B^{\prime}=\Omega$ holds, which yields three equations:
@@
The ITH method resolves this by exploiting heteroskedasticity. Suppose there are $S$ states such that the variance–covariance matrix $\Sigma$ changes across states, while $b_{r}$ and $b_{f}$ in $B$ remain constant. For each state $s=1,\dots,S$, Eqs. (7)–(9) hold, with $\omega_{r,s}$ and $\omega_{f,s}$ varying by state.
@@
the parameters are identified. [^8] See Rigobon (2003) for details on estimation and identification in general settings. Under these conditions, the parameters can be estimated by the generalized method of moments (GMM) using the $3S$ moment conditions implied by Eqs. (7)–(9).
@@ 4.3 SVAR Model
where $B$ and $\Omega$ are as in Eq. (5) and $\Phi_{j}$ are $2\times 2$ matrices capturing autocorrelation.
@@ 5.1 Macroeconomic News Effects
Figure 2: Average values of structural parameters ($b_{r}$, $b_{f}$, $\omega_{r}$, and $\omega_{f}$) and market activity variables ($D^{-1}$ and $SPR$) on days with macroeconomic announcements released at 9:00 (Chicago time; black circles, solid line) and on days without announcements (gray triangles, dashed line). Estimates of $b_{r}$ and $b_{f}$ are averaged over 15-minute intervals, whereas the other variables are averaged over five-minute intervals.
To formalize these patterns, we estimate regressions with two announcement dummies. The first, $ANN_{t}$, equals one if any of the eight announcements are released in interval $t$; the second, $ANN_{t}^{-}$, equals one if the announced value falls below the consensus forecast. Leads and lags of these dummies are included to capture pre- and post-announcement effects. Regressions are run separately for parameters computed over 15-minute intervals ($b_{r}$ and $b_{f}$) and over 5-minute intervals ($\omega_{r}$ and $\omega_{f}$), with controls for market activity variables—the reciprocal of depth $D^{-1}$ defined in Eq. (1), the number of order book events $NE$, the average size of events $ASE$, and the average spread SPRSPR—as well as time dummies.
@@
Table 2: Regression results with announcement dummies
@@
  • Notes: The dependent variables are the price impact ($b_{r}$), flow impact ($b_{f}$), return innovation volatility ($\omega_{r}$), and flow innovation volatility ($\omega_{f}$). Reported coefficients correspond to the announcement dummy ($ANN_{t}$), the negative-announcement dummy ($ANN_{t}^{-}$), and their leads and lags, when significant. Coefficients for other control variables are nearly identical to those in Table 5 and are omitted for brevity. $\bar{R}^{2}$ denotes the adjusted R-squared. Robust standard errors clustered by date (1,490 clusters) are reported in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
@@ 5.2 Structural Parameters
The reduced-form VAR model (12) is estimated using returns $r_{t}$ and order flow imbalances $f_{t}$ over one-second intervals, separately for each 15-minute interval and each day. This results in 38,740 estimations. [^10] The number of lags $p$ chosen by the Akaike information criterion is positive in 38,725 intervals. On average, lagged returns explain 8.7% of the variation in returns and lagged flows explain 4.6% of the variation in flows, with $R^{2}$ values occasionally exceeding 20%. These findings confirm strong auto- and cross-correlation structures, motivating the SVAR framework. The variance–covariance matrices of the residuals are then used to check the rank condition in Eq. (10), leaving 37,029 intervals suitable for ITH estimation.
For each valid interval, the ITH method is applied to estimate the structural parameters in Eq. (5). The price and flow impacts ($b_{r},b_{f}$) are assumed constant within a 15-minute window, while the standard deviations of return and flow innovations ($\omega_{r},\omega_{f}$) vary across three nested five-minute subintervals. This yields one set of $(b_{r},b_{f})$ and three sets of $(\omega_{r},\omega_{f})$ for each 15-minute block. Because three heteroskedastic states are used, the system is over-identified.
@@
Table 3: Estimation results for structural parameters
@@
  • Notes: For each 15-minute interval satisfying the rank condition in Eq. (10), the price and flow impacts ($b_{r}$ and $b_{f}$) and three sets of innovation volatilities ($\omega_{r,1}$, $\omega_{r,2}$, $\omega_{r,3}$, $\omega_{f,1}$, $\omega_{f,2}$, and $\omega_{f,3}$) are estimated using the ITH method, yielding 37,029 estimates in total. The last column (*) reports the proportion of intervals in which the coefficient is statistically significant, i.e., the absolute $t$-value exceeds two.
@@ 5.3 Impulse Responses
Return-to-Return Impact ($IRF_{rr})$ Flow-to-Return Impact ($IRF_{rf})$
@@
Return-to-Flow Impact ($IRF_{fr})$ Flow-to-Flow Impact ($IRF_{ff})$
@@
Figure 3: Impulse responses to a one-unit shock in return and flow innovations, up to 10 lags. $IRF_{ij}$ denotes the response of variable $i$ to a shock in variable $j$. The figure shows the median of the 37,029 estimated impulse responses. The shaded area indicates the 5th and 95th percentiles.
Table 4 reports summary statistics of the long-run impacts $I_{ij}(\infty)$ defined in Eq. (13). All are positive on average except for $I_{fr}(\infty)$ in some intervals. A one-unit flow shock, corresponding to an unexpected increase of about 1,000 contracts at the best bid or ask, raises prices by roughly 1.42 basis points in the long run. Using the mean flow volatility $\omega_{f}$ (0.34, about half the mean depth in Table 1), this translates into a price impact of 0.48 basis points—closely aligned with the 0.5–0.6 basis points reported by Hautsch and Huang (2012) for limit orders of half-depth.
@@
Table 4: Summary statistics of long-run impacts
@@
  • Notes: Summary statistics are based on 37,029 estimates of the long-run impacts $I_{ij}(\infty)$ defined in Eq. (13).
@@ 5.4 Intraday Variations
Figure 4: Intraday averages of estimated price and flow impacts ($b_{r}$ and $b_{f}$) and standard deviations of return and flow innovations ($\omega_{r}$ and $\omega_{f}$). Averages of $b_{r}$ and $b_{f}$ are computed for each 15-minute interval, whereas those of $\omega_{r}$ and $\omega_{f}$ are computed for each five-minute interval. Dashed lines indicate the full-sample averages.
@@
Table 5: Regression results for intraday variations
@@
  • Notes: The dependent variables are structural parameters ($b_{r}$ and $b_{f}$) estimated over 15-minute intervals, and innovation volatilities ($\omega_{r}$ and $\omega_{f}$) estimated over five-minute intervals. The explanatory variables include reciprocal depth $D^{-1}$ (per thousand), number of events $NE$ (in millions), average event size $ASE$ (in thousands), and average spread $SPR$, along with time dummies. Robust standard errors clustered by date (1,490 clusters) are reported in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.